# Global Continuation of Stable Periodic Orbits in Systems of Competing Predators

### v1

@@ 2.3 Verifying the contraction
$A_{\textnormal{tail}}\chi:=(0,0,0,u^{\prime}),\qquad(u_{j}^{\prime})_{k}:=\begin{cases}0,&|k|\leq K,\\ \displaystyle\frac{1}{ik}(u_{j})_{k},&|k|>K,\end{cases}\quad j=1,2,3.$ (41)
We emphasize that computing $A_{\textnormal{finite}}$ is done without numerically inverting the entire square matrix representing $\Pi_{N,K}DF(\bar{\chi})\Pi_{N,K}$. Since $DF(\bar{\chi})$ amounts to a multiplication operator with respect to $\eta$, we compute numerically $A_{\ell}\in\mathscr{B}(\Pi_{0,K}\mathcal{X}_{\nu},\Pi_{0,K}\mathcal{X}_{\nu})$ such that $A_{\ell}\approx(\Pi_{0,K}DF|_{\eta_{\ell}}(\bar{\chi}_{\ell})\Pi_{0,K})^{-1}$ and, as done for $\bar{\chi}$, we use the inverse discrete Fourier transform to obtain a Chebyshev polynomial associated to each component of $A_{\ell}$.
Recalling the domain of $F$ given in (37) and by the definition of $A_{\textnormal{tail}}$, it follows that
$T\in C^{2}\left(\left\{\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\mathcal{X}_{\nu}\,:\,\|(u_{3}+\alpha_{1})^{-1}\|_{\nu},\,\|(u_{3}+\alpha_{2})^{-1}\|_{\nu}<\infty\right\}\right).$ (42)
We denote by $B_{r}(\chi)$ the open ball in $\mathcal{X}_{\nu}$ with radius $r>0$ and centered at $\chi$, and by $\textnormal{cl}(B_{r}(\chi))$ its closure in $\mathcal{X}_{\nu}$. The next two lemmas provide sufficient conditions for $T$ to be, respectively, a self-mapping and a contraction which, by construction of $A$, is enough to obtain a zero of $F$.
Lemma 2.4 (Self-map).
Let $R>0$ and $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$. If there exists an $r\in(0,R]$ such that
$\|AF(\bar{\chi})\|_{\mathcal{X}_{\nu}}+r\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}+\frac{r^{2}}{2}\sup_{\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}\leq r,$ (43)
then $T:\textnormal{cl}(B_{r}(\bar{\chi}))\to\textnormal{cl}(B_{r}(\bar{\chi}))$.
Proof.
Let $\chi\in\textnormal{cl}(B_{r}(\bar{\chi}))$. Taylor’s theorem yields
$\|T(\chi)-\bar{\chi}\|_{\mathcal{X}_{\nu}}$ $=\|T(\bar{\chi})-\bar{\chi}+[DT(\bar{\chi})](\chi-\bar{\chi})+\int_{0}^{1}[DT(\bar{\chi}+t(\chi-\bar{\chi}))-DT(\bar{\chi})](\chi-\bar{\chi})\,\mathrm{d}t\|_{\mathcal{X}_{\nu}}$
$\leq\|AF(\bar{\chi})\|_{\mathcal{X}_{\nu}}+r\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}+\frac{r^{2}}{2}\sup_{\chi^{\prime}\in\textnormal{cl}(B_{r}(\bar{\chi}))}\|AD^{2}F(\chi^{\prime})\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$
$\leq r.$
Lemma 2.5 (Contraction).
Under the assumption of Lemma 2.4, if
$\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}+r\sup_{\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}<1,$ (44)
then $T$ is a contraction in $\textnormal{cl}(B_{r}(\bar{\chi}))$.
Proof.
Let $\chi,\chi^{\prime}\in\textnormal{cl}(B_{r}(\bar{\chi}))$. By the mean value theorem, we have
$\|T(\chi)-T(\chi^{\prime})\|_{\mathcal{X}_{\nu}}$ $\leq\sup_{\chi^{\prime\prime}\in\textnormal{cl}(B_{r}(\bar{\chi}))}\|DT(\bar{\chi})-DT(\bar{\chi})+DT(\chi^{\prime\prime})\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}\|\chi-\chi^{\prime}\|_{\mathcal{X}_{\nu}}$
$\leq\left(\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}+r\sup_{\chi^{\prime\prime}\in\textnormal{cl}(B_{r}(\bar{\chi}))}\|AD^{2}F(\chi^{\prime\prime})\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}\right)\|\chi-\chi^{\prime}\|_{\mathcal{X}_{\nu}}.$
Corollary 2.6.
$F$ has a unique zero in $\textnormal{cl}(B_{r}(\bar{\chi}))$.
Proof.
Since $\|ADF(\bar{\chi})-I\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}<1$, it follows that $ADF(\bar{\chi})$ is invertible, implying that $A$ is surjective. By definition, $A_{\textnormal{tail}}:\Pi_{>K}\mathcal{X}_{\nu}\to\Pi_{>K}\mathcal{X}_{\nu}$ is injective. By the construction of $A$, since $A$ is surjective, $A_{\textnormal{finite}}$ is also surjective. Since $A_{\textnormal{finite}}$ acts as a finite-dimensional square matrix, it is in fact injective. Therefore, the unique fixed-point of $T$ in $\textnormal{cl}(B_{r}(\bar{\chi}))$ is indeed a zero of $F$. ∎
@@
In the next sections, we derive practical estimates to verify the assumptions of Lemma 2.4 and Lemma 2.5. Specifically, we obtain formulas, in terms of the numerical approximation $\bar{\chi}$, to bound each norm involved in the inequalities (43)–(44), requiring only a finite, albeit large, number of calculations to be carried out by the computer.
@@
$W_{0}:=\begin{pmatrix}\rho(\bar{u})\\ \partial_{t}\bar{u}-\omega_{0}\end{pmatrix},$ (45)
@@
$[\mathcal{M}_{\phi}\psi](t,\eta):=\phi(t,\eta)\psi(\eta),\qquad\text{for all }\psi\in\mathcal{P}_{\nu}.$ (51)
For simplicity, we use the same symbol $\mathcal{M}_{\phi}$ for both operators. This slight abuse of notation should not cause confusion since the operators are essentially the same.
@@
$W_{1}:=\begin{pmatrix}0&0&0&\rho\\ -\omega_{2}&-\omega_{3}&-\omega_{4}&\partial_{t}-\omega_{1}\end{pmatrix},$ (52)
@@
$\omega_{1}\phi$ $:=\sum_{j=1}^{3}\begin{pmatrix}\omega_{1}^{(1,j)}\phi_{j}\\ \omega_{1}^{(2,j)}\phi_{j}\\ \omega_{1}^{(3,j)}\phi_{j}\end{pmatrix},\qquad$ for all $\text{for all }\phi=(\phi_{1},\phi_{2},\phi_{3})\in\mathcal{U}_{\nu},$ (53)
$\omega_{j}\psi$ $:=\begin{pmatrix}\omega_{j}^{(1)}\psi\\ \omega_{j}^{(2)}\psi\\ \omega_{j}^{(3)}\psi\end{pmatrix},\qquad$ for $\text{for all }\psi\in\mathcal{P}_{\nu},\quad j=2,3,4,$ (54)
@@
2.3.3 Estimate for $\sup_{\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$
@@
$\sup_{\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}\leq\|A\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}\sup_{\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|D^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))},$
it remains to bound $\|D^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$ for all $\chi\in\textnormal{cl}(B_{R}(\bar{\chi}))$. Introducing the notation $\nabla=(\partial_{\tau},\partial_{\zeta_{1}},\partial_{\zeta_{2}},\partial_{u_{1}},\partial_{u_{2}},\partial_{u_{3}})$, and to simplify the notation $f_{j}=f_{j}(\zeta_{1},\zeta_{2},u)$, we have
@@
To compute the above norms, we need to bound $(u_{3}+\alpha_{j})^{-p}$ for $p=1,2,3$ and $u_{3}\in\textnormal{cl}(B_{R}(\bar{u}_{3}))$. The following lemma provides a means for $p=1$, while the cases $p=2,3$ are handled by using the Banach algebra properties given in (19) and (25).
@@
$\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\|\phi^{-1}\|_{\nu}\leq\frac{\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}}{1-\|1-\bar{\phi}\bar{\phi}_{\textnormal{inv}}\|_{\nu}-R\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}}.$ (61)
@@
$\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\|\phi^{-1}\|_{\nu}$ $=\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\|\bar{\phi}_{\textnormal{inv}}(\phi\bar{\phi}_{\textnormal{inv}})^{-1}\|_{\nu}$
$=\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\|\bar{\phi}_{\textnormal{inv}}(1-(1-\phi\bar{\phi}_{\textnormal{inv}}))^{-1}\|_{\nu}$
$=\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\|\bar{\phi}_{\textnormal{inv}}\sum_{n\geq 0}(1-\phi\bar{\phi}_{\textnormal{inv}})^{n}\|_{\nu}$
$\leq\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}\sup_{\phi\in\textnormal{cl}(B_{R}(\bar{\phi}))}\sum_{n\geq 0}\|1-\phi\bar{\phi}_{\textnormal{inv}}\|_{\nu}^{n}$
@@
$\sup_{\chi=(\tau,\zeta_{1},\zeta_{2},u)\in\textnormal{cl}(B_{R}(\bar{\chi}))}\|\partial_{\tau}\nabla(\tau f_{j})\|$ $=\max\left(\|\delta_{j}\frac{u_{3}-\lambda_{j}}{u_{3}+\alpha_{j}}\|_{\nu},\|\delta_{j}\frac{\lambda_{j}+\alpha_{j}}{(u_{3}+\alpha_{j})^{2}}u_{j}\|_{\nu}\right)$
@@ 2.4 Real-valued solutions and boundary crossing
Given $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, if Corollary 2.6 holds, then there exist $r>0$, and a unique $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in\textnormal{cl}(B_{r}(\bar{\chi}))$ satisfying $F(\tilde{\chi})=0$. Whence, from (7), we have
@@
$\Sigma(\chi):=(\tau^{*},\zeta_{1}^{*},\zeta_{2}^{*},\Sigma_{0}(u_{1}),\Sigma_{0}(u_{2}),\Sigma_{0}(u_{3})),$ (64)
@@
$(\Sigma_{0}(\phi))_{n,k}:=\phi_{n,-k}^{*},\qquad n\in\mathbb{N}_{0},\quad k\in\mathbb{Z}.$ (66)
@@
Suppose that Corollary 2.6 holds, that is, there exist $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, $r>0$, and a unique $\tilde{\chi}\in\textnormal{cl}(B_{r}(\bar{\chi}))$ such that $F(\tilde{\chi})=0$. If $\Sigma(\bar{\chi})=\bar{\chi}$, then $\Sigma(\tilde{\chi})=\tilde{\chi}$.
@@
Let $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, $r>0$ and $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in\textnormal{cl}(B_{r}(\bar{\chi}))$. For any $\eta\in[-1,1]$, we have
@@
The previous lemma can only provide an enclosure of the exact value. So the sign of $\tilde{\zeta}_{j}$ becomes undetermined as we approach $\hat{\eta}_{j}$, assuming there exists $\hat{\eta}_{1},\hat{\eta}_{2}\in(-1,1)$ such that $\tilde{\zeta}_{2}(\hat{\eta}_{1})=0$ and $\tilde{\zeta}_{1}(\hat{\eta}_{2})=0$. Let $h_{1}^{-},h_{1}^{+},h_{2}^{-}$, and $h_{2}^{+}$ satisfy $-1<h_{1}^{-}<h_{1}^{+}\leq 0\leq h_{2}^{-}<h_{2}^{+}<1$. Hence, to conclude that the family crosses the boundary plane $Q_{1}$ (resp., $Q_{2}$) only at some $\hat{\eta}_{1}\in(h_{1}^{-},h_{1}^{+})$ (resp., $\hat{\eta}_{2}\in(h_{2}^{-},h_{2}^{+})$), we need to show:
@@
Let $\bar{\psi}\in\Pi_{N}\mathcal{P}_{\nu}$, $\nu>1$, $r>0$ and $\tilde{\psi}\in\textnormal{cl}(B_{r}(\bar{\psi}))$. It follows that

### v2

@@ 2.2 Newton–Kantorovich type theorem
$A_{\textnormal{tail}}\chi\coloneqq(0,0,0,u^{\prime}),\qquad(u_{j}^{\prime})_{k}\coloneqq\begin{cases}0,&|k|\leq K,\\ \displaystyle\frac{1}{ik}(u_{j})_{k},&|k|>K,\end{cases}\quad j=1,2,3.$ (50)
On the other hand, $A_{\textnormal{finite}}\in\mathscr{B}(\Pi_{K}\mathcal{X}_{\nu},\Pi_{K}\mathcal{X}_{\nu})$ is an approximation of the inverse $(\Pi_{K}DF(\bar{\chi})\Pi_{K})^{-1}$ constructed by interpolating each entries of the $N+1$ matrices $(\Pi_{0,K}DF|_{\eta_{\ell}}(\bar{\chi})\Pi_{0,K})^{-1}$ with a Chebyshev polynomial.
@@
In the next sections, we derive practical estimates to verify the assumptions of Theorem 2.1. Specifically, we obtain explicit bounds for each norm in (45), expressed in terms of the numerical approximation $\bar{\chi}$ and so that only a finite, potentially substantial, set of computer calculations is needed.
@@
$W_{0}\coloneqq\begin{pmatrix}\rho(\bar{u})\\ \partial_{t}\bar{u}-\omega_{0}\end{pmatrix}$, (51)
@@
[…]
For simplicity, we use the same symbol $\mathcal{M}_{\phi}$ for both operators. This slight abuse of notation should not cause confusion since the operators are essentially of the same nature.
@@
$W_{1}\coloneqq\begin{pmatrix}0&0&0&\rho\\ -\omega_{2}&-\omega_{3}&-\omega_{4}&\partial_{t}-\omega_{1}\end{pmatrix}$, (58)
@@
$\omega_{1}\phi$ $\coloneqq\sum_{j=1}^{3}\begin{pmatrix}\omega_{1}^{(1,j)}\phi_{j}\\ \omega_{1}^{(2,j)}\phi_{j}\\ \omega_{1}^{(3,j)}\phi_{j}\end{pmatrix}$$\qquad$ $\text{for all }\phi=(\phi_{1},\phi_{2},\phi_{3})\in\mathcal{U}_{\nu},$ (59)
$\omega_{j}\psi$ $\coloneqq\begin{pmatrix}\omega_{j}^{(1)}\psi\\ \omega_{j}^{(2)}\psi\\ \omega_{j}^{(3)}\psi\end{pmatrix}$$\qquad$ $\text{for all }\psi\in\mathcal{P}_{\nu},\quad j=2,3,4,$ (60)
@@
2.2.3 Estimate for $\sup_{\chi\in B_{R}(\bar{\chi})}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$
@@
$\sup_{\chi\in B_{R}(\bar{\chi})}\|AD^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}\leq\|A\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu})}\sup_{\chi\in B_{R}(\bar{\chi})}\|D^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))},$
it remains to bound $\|D^{2}F(\chi)\|_{\mathscr{B}(\mathcal{X}_{\nu},\mathscr{B}(\mathcal{X}_{\nu},\mathcal{X}_{\nu}))}$ for all $\chi\in B_{R}(\bar{\chi})$. Introducing the notation $\nabla=(\partial_{\tau},\partial_{\zeta_{1}},\partial_{\zeta_{2}},\partial_{u_{1}},\partial_{u_{2}},\partial_{u_{3}})$, and to simplify the notation $f_{j}=f_{j}(\zeta_{1},\zeta_{2},u)$, we have
@@
To compute the above norms, we need to bound $(u_{3}+\alpha_{j})^{-p}$ for $p=1,2,3$ and $u_{3}\in B_{R}(\bar{u}_{3})$. The following lemma provides a means for $p=1$, while the cases $p=2,3$ are handled by using the Banach algebra properties given in (32) and (34).
@@
$\sup_{\phi\in(B_{R}(\bar{\phi})}\|\phi^{-1}\|_{\nu}\leq\frac{\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}}{1-\|1-\bar{\phi}\bar{\phi}_{\textnormal{inv}}\|_{\nu}-R\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}}.$ (67)
@@
$\sup_{\phi\in B_{R}(\bar{\phi})}\|\phi^{-1}\|_{\nu}$ $=\sup_{\phi\in B_{R}(\bar{\phi})}\|\bar{\phi}_{\textnormal{inv}}(\phi\bar{\phi}_{\textnormal{inv}})^{-1}\|_{\nu}$
$=\sup_{\phi\in B_{R}(\bar{\phi})}\|\bar{\phi}_{\textnormal{inv}}(1-(1-\phi\bar{\phi}_{\textnormal{inv}}))^{-1}\|_{\nu}$
$=\sup_{\phi\in B_{R}(\bar{\phi})}\|\bar{\phi}_{\textnormal{inv}}\sum_{n\geq 0}(1-\phi\bar{\phi}_{\textnormal{inv}})^{n}\|_{\nu}$
$\leq\|\bar{\phi}_{\textnormal{inv}}\|_{\nu}\sup_{\phi\in B_{R}(\bar{\phi})}\sum_{n\geq 0}\|1-\phi\bar{\phi}_{\textnormal{inv}}\|_{\nu}^{n}$
@@
$\sup_{\chi=(\tau,\zeta_{1},\zeta_{2},u)\in B_{R}(\bar{\chi})}\|\partial_{\tau}\nabla(\tau f_{j})\|$ $=\max\left(\|\delta_{j}\frac{u_{3}-\lambda_{j}}{u_{3}+\alpha_{j}}\|_{\nu},\|\delta_{j}\frac{\lambda_{j}+\alpha_{j}}{(u_{3}+\alpha_{j})^{2}}u_{j}\|_{\nu}\right)$
@@ 2.3 Real-valued solutions and boundary crossing
Given $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, if Theorem 2.1 holds, then there exists a unique $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in B_{r}(\bar{\chi})$, for some $r>0$, satisfying $F(\tilde{\chi})=0$. Whence, from (18), we have
@@
$\Sigma(\chi)\coloneqq(\tau^{*},\zeta_{1}^{*},\zeta_{2}^{*},\Sigma_{0}(u_{1}),\Sigma_{0}(u_{2}),\Sigma_{0}(u_{3})),$ (70)
@@
$(\Sigma_{0}(\phi))_{n,k}\coloneqq\phi_{n,-k}^{*},\qquad n\in\mathbb{N}_{0},\quad k\in\mathbb{Z}.$ (72)
@@
Suppose that Theorem 2.1 holds, that is, there exist $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, and a unique $\tilde{\chi}\in B_{r}(\bar{\chi})$, for some $r>0$, such that $F(\tilde{\chi})=0$. If $\Sigma(\bar{\chi})=\bar{\chi}$, then $\Sigma(\tilde{\chi})=\tilde{\chi}$.
@@
Let $\bar{\chi}\in\Pi_{N,K}\mathcal{X}_{\nu}$, $r>0$ and $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in B_{r}(\bar{\chi})$. For any $\eta\in[-1,1]$, we have
@@
The previous lemma can only provide an enclosure of the exact value. So the sign of $\tilde{\zeta}_{j}$ becomes undetermined as we approach $\hat{\eta}_{j}$, assuming there exists $\hat{\eta}_{1},\hat{\eta}_{2}\in(-1,1)$ such that $\tilde{\zeta}_{2}(\hat{\eta}_{1})=0$ and $\tilde{\zeta}_{1}(\hat{\eta}_{2})=0$. Let $h_{1}^{-},h_{1}^{+},h_{2}^{-}$, and $h_{2}^{+}$ satisfy $1<h_{1}^{-}<h_{1}^{+}\leq 0\leq h_{2}^{-}<h_{2}^{+}<1$. Hence, to conclude that the family crosses the boundary plane $Q_{1}$ (resp., $Q_{2}$) only at some $\hat{\eta}_{1}\in(h_{1}^{-},h_{1}^{+})$ (resp., $\hat{\eta}_{2}\in(h_{2}^{-},h_{2}^{+})$), we need to show:
@@
Let $\bar{\psi}\in\Pi_{N}\mathcal{P}_{\nu}$, $\nu>1$, $r>0$ and $\tilde{\psi}\in B_{r}(\bar{\psi})$. It follows that